Determine the equation of the line passing through the point (2,9)and perpendicular to the line with equation 5x - y =7

Answer by mananth(16949)

5x-y=7
Find the slope of this line
-y=-5x-7
Divide by -1
y =5x+ 7
Compare this equation with y=mx+b
slope m =5.00

The slope of a line perpendicular to the above line will be the negative reciprocal
m1*m2=-1
The slope of the required line will be -0.20 -0.2 =

m=-0.20 ,point (2,9)
Find b by plugging the values of m & the point in
y=mx+b
9=-0.4+b
b=9.4
m=-0.2
The required equation is y=-0.2x+ 9.4